A mother is 33 years older than her twin children. In 3 years her age will be twice the combined ages of the twins. How old are the twins now?
Answer by ankor@dixie-net.com(22740)

let t = the age of the twins
:
A mother is 33 years older than her twin children.
m = t + 33
In 3 years her age will be twice the combined ages of the twins.
m + 3 = 2(2(t+3))
m + 3 = 4t + 12
m = 4t + 12 - 3
m = 4t + 9
How old are the twins now?
Replace m with t + 33
t + 33 = 4t + 9
33 - 9 = 4t - t
24 = 3t
t = 24/3
t = 8 yrs old the twins now
:
:
Check this, find Mom's age
m = 8 + 33
m = 41 yrs old
:
Check in the statement
"In 3 years her age will be twice the combined ages of the twins."
41 + 3 = 2(2(8+3))
44 = 2(22)